OpenStudy (anonymous):

What is the solution to 1/6a -1/3=2/3

OpenStudy (anonymous):

a=1/6

OpenStudy (anonymous):

For such questions, think of a way of rearranging the problem statement so you have all the terms with variables on one side, and rest on the other side. So rearranging would give us 1/6a = 2/3 + 1/3 => 1/6a = 3/3 => 1/6a = 1 now multiply both sides of the equation with 'a' you will get a = 1/6

OpenStudy (anonymous):

this is not an answer in my book

OpenStudy (helder_edwin):

is the a multiplying 1/6 or only the 6?

OpenStudy (anonymous):

What answer does your book give? Maybe you made an error in the way you wrote the question? Could you check? :)

OpenStudy (helder_edwin):

i mean is it \[\frac{1}{6} a\] ?

OpenStudy (anonymous):

no i wrote it correct here are what the answers could be -2 or 6 or -3/4 or 1/4

OpenStudy (helder_edwin):

so the answer is 6. and you wrote the question improperly.

OpenStudy (anonymous):

no it is correct

OpenStudy (helder_edwin):

you should have writen (1/6)a-1/3=2/3

OpenStudy (helder_edwin):

this way you have \[\frac{1}{6} 6 - \frac{1}{3}=\frac{2}{3}\]